Note that an item can be classified as good (G), ok (O), and not good (N). Consider randomly selected two items. a) Define an event A that represents both items have the same classification. b) What is the probability of the above event?
Answer #1

a) Let the event GN indicate that the first item is good and second item is not good. Then,

A = {GG, OO, NN}

b) Number of cases in sample space = Number of classifications possible for first item x Number of classifications possible for second item

= 3 x 3

= 9

So, P(A) = Number of cases in event A / Total number of cases possible

= 3/(3x3)

= 3/9

= 1/3